Betty and Bob own a coupon bond priced at $100 with a Macaulay Duration of 10.4 years and a convexity of 100 as per annum. If rates decrease from 8% per annum compounded semiannually to 7% per annum compounded semiannually then find the approximate new price of the bond, correct to 3 places after the decimal point.
Answer #1

Modified Duration = 10.40/1.08 = 9.6296 years

Change in Price = -Mod D(Change in YTM) + ).50(C)(Change in YTM)2

Change in Price = -9.6296(-0.01) + 0.50(100)(-0.01)2

Change in Price = 10.1296%

New Price = (1 + 0.101296)(100)

New Price = $110.1296